Given the Equation 6K + B2O3. 3K2O + 2B


How many grams of K2O will be produced whe you react one 3.5 g of B2O3

1
Expert's answer
2022-05-31T15:31:04-0400

Solution:

The molar mass of B2O3 is 69.63 g/mol

Therefore,

Moles of B2O3 = (3.5 g B2O3) × (1 mol B2O3 / 69.63 g B2O3) = 0.05026 mol B2O3


Balanced chemical equation:

6K + B2O3 → 3K2O + 2B

According to stoichiometry:

1 mol of B2O3 reacts with 3 mol of K2O

Thus, 0.05026 mol of B2O3 react with:

(0.05026 mol B2O3) × (3 mol K2O / 1 mol B2O3) = 0.15078 mol K2O


The molar mass of K2O is 94.2 g/mol

Therefore,

Mass of K2O = (0.15078 mol K2O) × (94.2 g K2O / 1 mol K2O) = 14.2035 g K2O = 14.2 g K2O

Mass of K2O = 14.2 g


Answer: 14.2 grams of K2O will be produced